OpenStudy (anonymous):

please dont solve help on the first step: Let f(x) = x2 – 81. Find f–1(x).

OpenStudy (anonymous):

y = x^2 - 81 x = y^2 - 81 plus/minus sqrt(x + 81) = y

OpenStudy (anonymous):

\[ f^{-1} = \pm \sqrt{x+81}\]

OpenStudy (anonymous):

the answer is: f^1=+-√x+81

OpenStudy (anonymous):